OpenStudy (anonymous):

how do you do elimination using multiplication ? example problem : x-y=-8 7x+5y=16

hartnn (hartnn):

multiply the equation x-y=-8 with 5 what u get ?

hero (hero):

0. Given x - y = -8 7x+5y = 16 1. Multiply the first equation by 5 to get: 5x - 5y = -40 7x +5y = 16 2. Add both equations together to get: 12x = -24 3. Finish solving for x.

hero (hero):

If you already found x, you still have to find y.

OpenStudy (anonymous):

I Did Already

hero (hero):

Well, you're supposed to tell me what you found